# Cooperation of Experts: Fusing Heterogeneous Information with Large Margin

### v1

@@ 1 Introduction
Most real-world objects and data are heterogeneous in the form of multiple types or diverse forms of interactions, which pose significant challenges for modeling their intricate relationships (Han, 2009; shen2025heterophily). For instance, multimodal data combines diverse sources, such as images and text, where each modality contributes unique features that collectively offer a richer understanding of the underlying phenomena; in a social network, people are linked with different types of ties: friendship, family relationship, professional relationship, etc. To characterize the complexities of this heterogeneity, we encode multi-typed information into a unified heterogeneous multiplex network, where each layer contains the same number of nodes (possibly with different attributes) but a different type of links, enabling a comprehensive and effective representation of the diverse relationships inherent in real-world data (Sun & Han, 2013).
Inspired by the success of graph neural networks (GNNs) (Kipf & Welling, 2016; Wang et al., 2019), numerous approaches have been proposed for multiplex network representation learning (Jing et al., 2021; Shen et al., 2024b). However, these methods often rely on training a single predictor across the entire network, which can lead to suboptimal results by neglecting the intrinsic heterogeneity and varying characteristics of different link patterns (Wang et al., 2024). As shown in Figures 1(a) and 1(b), classifier performance, which is independently trained on each network, differs significantly across different layers, highlighting the importance of accounting for this variability. To effectively capture diverse node patterns, dedicated experts—components designed to focus on specific aspects of the network, are introduced (Shi et al., 2024b). Each expert specializes in learning from a particular type of interaction, thereby reducing cross-interference between relationships and enabling a more nuanced representation of complex structures.
When utilizing multiple experts, effective collaboration among them becomes crucial. Existing Mixture of Experts (MoE) approaches employ a gating mechanism that activates only a subset of experts during both training and inference (Li et al., 2023a; Shi et al., 2024a). While this strategy improves efficiency, it inherently limits the ability to fully exploit the rich and diverse information embedded in heterogeneous data. This highlights the importance of fostering collaboration rather than competition among experts to leverage cross-layer knowledge effectively. However, this shift introduces two key challenges that must be addressed.
@@ 2.1 Multiplex Network Learning
Multiplex network learning focuses on capturing diverse structural patterns in multi-relational networks to support tasks such as node classification, clustering, and link prediction (qian2024upper; xie2025one). With the rise of GNNs, HAN (Wang et al., 2019) integrates structural and attribute information using attention mechanisms at both the node and semantic levels, while MAGNN (Fu et al., 2020) and MHGCN (Yu et al., 2022) leverage graph convolution to extract relational features across multiple layers. MGBO (Huang et al., 2024) employs self-expression learning for homophily modeling, whereas HDMI (Jing et al., 2021) and DMG (Mo et al., 2023) utilize contrastive learning to enhance self-supervised representations. However, these methods heavily rely on reliable network topology structures, limiting their effectiveness in real-world scenarios with noisy or incomplete data.
@@ 2.2 Graph Structure Learning
Graph Structure Learning (GSL) has become a crucial area of research in GNNs, particularly for optimizing unreliable graph structures with significant noise. Supervised GSL methods often generate new adjacency matrices with learnable components, optimized using label information. These methods include probabilistic models (e.g., LDS and GEN (Franceschi et al., 2019; Wang et al., 2021)), similarity-based techniques (e.g., GRCN and IDGL (Yu et al., 2021; Chen et al., 2020)), attention mechanisms (e.g., NodeFormer (Wu et al., 2022)), or by treating all elements in the adjacency matrix as learnable parameters (e.g., ProGNN (Jin et al., 2020)). In contrast, methods like SUBLIME, STABLE, and GSR (Liu et al., 2022; Li et al., 2022; Zhao et al., 2023) leverage contrastive learning to construct graph structures in a self-supervised manner. While most previous studies have focused on single, homogeneous graphs, recent efforts have expanded GSL to heterogeneous graphs in supervised settings (li2024pc; pan2023beyond), as demonstrated by GTN and HGSL (Yun et al., 2019; Zhao et al., 2021). More recently, InfoMGF (Shen et al., 2024b) has explored unsupervised fusion of graphs from multiple sources, advancing research in multiplex graph structure learning. Despite these strides, there remains a lack of research addressing the challenge of effectively enabling the integration of complex information across multiple networks.
@@ 2.3 Multiple Learners Learning
Another notable specialization within this domain is the expert mechanism. The classical Mixture of Experts (MoE) framework enhances adaptivity by assigning specialized submodels (i.e., experts) to distinct regions of the input space (Li et al., 2023a). A key limitation of many GNNs is their homogeneous processing of the entire graph, which overlooks structural and feature diversity. To address this, recent studies integrate MoE with graph learning (Wang et al., 2024; Kim et al., 2023) to improve adaptability. Furthermore, Mowst (Zeng et al., 2024) refines this approach by training experts to specialize in both feature and structure modalities. However, existing expert mechanisms primarily focus on expert competition, often neglecting the effective utilization of inter-relational dependencies—precisely the gap our work aims to bridge.
@@ 4.1 Encoding Heterogeneous Information
$H_{v}$ is subsequently utilized to reconstruct the network using the $K$-nearest neighbors ($K$NN) method. In practice, we employ $K$NN sparsification with its locality-sensitive approximation to enhance efficiency (Fatemi et al., 2021). Once the reconstructed adjacency matrix is obtained, we perform post-processing operations, like existing GSL paradigm (Li et al., 2023b), to ensure that the final adjacency matrix $A_{v}^{\prime}$ adheres to key properties, including non-negativity, symmetry, and normalization. This process ultimately results in the construction of the heterogeneous multiplex networks $G=\{G_{1}^{\prime},\cdots,G_{V}^{\prime}\}$, where $G_{i}^{\prime}=\{A_{i}^{\prime},X^{i}\}$.
@@ 4.2 Two-level Design for Task-driven Experts
As described above, the two-level expert framework is designed to operate across distinct knowledge fields. Specifically, low-level experts focus on individual networks, learning network-specific link patterns, while high-level experts capture shared information across networks to identify higher-order dependencies that cannot be uncovered when networks are analyzed independently. To enhance the task relevance of low-level experts, we guide their learning by formulating it as the maximization of $I(G_{i}^{\prime};Y)$. For high-level experts, we incorporate $I(G_{i}^{\prime};G_{j}^{\prime})$ as a maximization objective during the training phase, which has been theoretically shown to provide a lower bound for $I(G_{i}^{\prime};G_{j})+I(G_{i};G_{j}^{\prime})$ (Federici et al., 2020). The loss function can be formulated as:
@@
Given a network $G$ with label $Y$, the cross-entropy loss $\mathcal{L}_{cls}(Z,Y)$ is the upper bound of $-I(G;Y)$, where $Z$ denotes the node representations of all nodes in network $G$. (Sun et al., 2022; Li et al., 2024)
@@ Appendix B Hyper-parameters Settings and Infrastructure
Our model is trained using the Adam optimizer. The hyper-parameter settings for all datasets are presented in Table 5. Here, $E$ represents the number of training epochs, which is tuned within the set $\{100,200,400,500,800,1000\}$ and $lr$ is the learning rate which is searched within the set $\{0.0001,0.005,0.001,0.005,0.01\}$. The hidden layer dimension $d_{h}$ and the representation dimension $d$ of the graph encoder GCN are tuned within the set $\{32,64,128,256\}$. The number of neighbors $K$ for $K$NN is searched within the set $\{5,10,15,20,30\}$. The order of graph aggregation $r$ and the number of layers $L$ in GCN are set to either 2 or 3, which is in line with the common layer count of GNN models (Baranwal et al., 2022). The temperature parameter $\tau_{c}$ in the contrastive loss is fixed at 0.2. In the optimization phase, the parameters $\alpha$ and $\gamma$ are both set to 100. For the large-scale datasets MAG and Amazon, when computing the contrastive loss for estimating mutual information, we use a batch size of 2560 and process the data in batches.

### v2

@@ 1 Introduction
Most real-world objects and data are heterogeneous in the form of multiple types or diverse forms of interactions, which pose significant challenges for modeling their intricate relationships (Han, 2009; Shen & Kang, 2025). For instance, multimodal data combines diverse sources, such as images and text, where each modality contributes unique features that collectively offer a richer understanding of the underlying phenomena; in a social network, people are linked with different types of ties: friendship, family relationship, professional relationship, etc. To characterize the complexities of this heterogeneity, we encode multi-typed information into a unified heterogeneous multiplex network, where each layer contains the same number of nodes (possibly with different attributes) but a different type of links, enabling a comprehensive and effective representation of the diverse relationships inherent in real-world data (Sun & Han, 2013).
Inspired by the success of graph neural networks (GNNs) (Kipf & Welling, 2016; Wang et al., 2019), numerous approaches have been proposed for multiplex network representation learning (Jing et al., 2021; Shen et al., 2024b). However, these methods often rely on training a single predictor across the entire network, which can lead to suboptimal results by neglecting the intrinsic heterogeneity and varying characteristics of different link patterns (Wang et al., 2024). As shown in Figures 1(a) and 1(b), classifier performance, which is independently trained on each network, differs significantly across different layers, highlighting the importance of accounting for this variability. To effectively capture diverse node patterns, dedicated experts—components designed to focus on specific aspects of the network, are introduced (Shi et al., 2024). Each expert specializes in learning from a particular type of interaction, thereby reducing cross-interference between relationships and enabling a more nuanced representation of complex structures.
When utilizing multiple experts, effective collaboration among them becomes crucial. Existing Mixture of Experts (MoE) approaches employ a gating mechanism that activates only a subset of experts during both training and inference (Li et al., ; Shi et al., ). While this strategy improves efficiency, it inherently limits the ability to fully exploit the rich and diverse information embedded in heterogeneous data. This highlights the importance of fostering collaboration rather than competition among experts to leverage cross-layer knowledge effectively. However, this shift introduces two key challenges that must be addressed.
@@ 2.1 Multiplex Network Learning
Multiplex network learning focuses on capturing diverse structural patterns in multi-relational networks to support tasks such as node classification, clustering, and link prediction (Qian et al., 2024; Xie et al., 2025). With the rise of GNNs, HAN (Wang et al., 2019) integrates structural and attribute information using attention mechanisms at both the node and semantic levels, while MAGNN (Fu et al., 2020) and MHGCN (Yu et al., 2022) leverage graph convolution to extract relational features across multiple layers. MGBO (Huang et al., 2024) employs self-expression learning for homophily modeling, whereas HDMI (Jing et al., 2021) and DMG (Mo et al., 2023) utilize contrastive learning to enhance self-supervised representations. However, these methods heavily rely on reliable network topology structures, limiting their effectiveness in real-world scenarios with noisy or incomplete data.
@@ 2.2 Graph Structure Learning
Graph Structure Learning (GSL) has become a crucial area of research in GNNs, particularly for optimizing unreliable graph structures with significant noise. Supervised GSL methods often generate new adjacency matrices with learnable components, optimized using label information. These methods include probabilistic models (e.g., LDS and GEN (Franceschi et al., 2019; Wang et al., 2021)), similarity-based techniques (e.g., GRCN and IDGL (Yu et al., 2021; Chen et al., 2020)), attention mechanisms (e.g., NodeFormer (Wu et al., 2022)), or by treating all elements in the adjacency matrix as learnable parameters (e.g., ProGNN (Jin et al., 2020)). In contrast, methods like SUBLIME, STABLE, and GSR (Liu et al., 2022; Li et al., 2022; Zhao et al., 2023) leverage contrastive learning to construct graph structures in a self-supervised manner. While most previous studies have focused on single, homogeneous graphs, recent efforts have expanded GSL to heterogeneous graphs in supervised settings (Li et al., 2024a; Pan & Kang, 2023), as demonstrated by GTN and HGSL (Yun et al., 2019; Zhao et al., 2021). More recently, InfoMGF (Shen et al., 2024b) has explored unsupervised fusion of graphs from multiple sources, advancing research in multiplex graph structure learning. Despite these strides, there remains a lack of research addressing the challenge of effectively enabling the integration of complex information across multiple networks.
@@ 2.3 Multiple Learners Learning
Another notable specialization within this domain is the expert mechanism. The classical Mixture of Experts (MoE) framework enhances adaptivity by assigning specialized submodels (i.e., experts) to distinct regions of the input space (Li et al., ). A key limitation of many GNNs is their homogeneous processing of the entire graph, which overlooks structural and feature diversity. To address this, recent studies integrate MoE with graph learning (Wang et al., 2024; Kim et al., 2023) to improve adaptability. Furthermore, Mowst (Zeng et al., 2024) refines this approach by training experts to specialize in both feature and structure modalities. However, existing expert mechanisms primarily focus on expert competition, often neglecting the effective utilization of inter-relational dependencies—precisely the gap our work aims to bridge.
@@ 4.1 Encoding Heterogeneous Information
$H_{v}$ is subsequently utilized to reconstruct the network using the $K$-nearest neighbors ($K$NN) method. In practice, we employ $K$NN sparsification with its locality-sensitive approximation to enhance efficiency (Fatemi et al., 2021). Once the reconstructed adjacency matrix is obtained, we perform post-processing operations, like existing GSL paradigm (Li et al., 2023), to ensure that the final adjacency matrix $A_{v}^{\prime}$ adheres to key properties, including non-negativity, symmetry, and normalization. This process ultimately results in the construction of the heterogeneous multiplex networks $G=\{G_{1}^{\prime},\cdots,G_{V}^{\prime}\}$, where $G_{i}^{\prime}=\{A_{i}^{\prime},X^{i}\}$.
@@ 4.2 Two-level Design for Task-driven Experts
As described above, the two-level expert framework is designed to operate across distinct knowledge fields. Specifically, low-level experts focus on individual networks, learning network-specific link patterns, while high-level experts capture shared information across networks to identify higher-order dependencies that cannot be uncovered when networks are analyzed independently. To enhance the task relevance of low-level experts, we guide their learning by formulating it as the maximization of $I(G_{i}^{\prime};Y)$. For high-level experts, we incorporate $I(G_{i}^{\prime};G_{j}^{\prime})$ as a maximization objective during the training phase, which has been theoretically shown to provide a lower bound for $I(G_{i}^{\prime};G_{j})+I(G_{i};G_{j}^{\prime})$ (Federici et al., ). The loss function can be formulated as:
@@
Given a network $G$ with label $Y$, the cross-entropy loss $\mathcal{L}_{cls}(Z,Y)$ is the upper bound of $-I(G;Y)$, where $Z$ denotes the node representations of all nodes in network $G$. (Sun et al., 2022; Li et al., 2024b)
@@ Appendix B Hyper-parameters Settings and Infrastructure
Our model is trained using the Adam optimizer. The hyper-parameter settings for all datasets are presented in Table 5. Here, $E$ represents the number of training epochs, which is tuned within the set $\{100,200,400,500,800,1000\}$ and $lr$ is the learning rate which is searched within the set $\{0.0001,0.005,0.001,0.005,0.01\}$. The hidden layer dimension $d_{h}$ and the representation dimension $d$ of the graph encoder GCN are tuned within the set $\{32,64,128,256\}$. The number of neighbors $K$ for $K$NN is searched within the set $\{5,10,15,20,30\}$. The order of graph aggregation $r$ and the number of layers $L$ in GCN are set to either 2 or 3, which is in line with the common layer count of GNN models (Baranwal et al., ). The temperature parameter $\tau_{c}$ in the contrastive loss is fixed at 0.2. In the optimization phase, the parameters $\alpha$ and $\gamma$ are both set to 100. For the large-scale datasets MAG and Amazon, when computing the contrastive loss for estimating mutual information, we use a batch size of 2560 and process the data in batches.